AAA62 – Schedule of Lectures

The talks are scheduled for 20 minutes in length, followed by a 5 minute discussion period and a 5 minute break in order to allow participants to change sessions.

Friday, June 15th, Morning

Lecture Hall HS 3

Chair: Dragan Mašulović

Lecture Hall HS 5

Chair: Miroslav Ploščica

Lecture Hall HS 7

Chair: Reinhard Pöschel

9:00 - 10:00 Bruno Buchberger

Gröbner Bases and Theorema

10:15 - 10:45 Klaus Denecke

Presolid varieties of semirings

Helmut Länger

Restricted congruence regularity of algebras (joint work with Ivan Chajda)

Andrei Krokhin

Classifying the complexity in Allen’s algebra

10:45 - 11:15 Shelly L. Wismath

Short-solid superassociative varieties of type (n)

Günther Eigenthaler

Balanced congruences

Franz Binder

Computational Methods for Transformation Nearrings

11:15 - 11:45  Coffee break

11:45 - 12:15 Dietmar Schweigert

Equivalent hypersubstitutions and the Mal’cev correspondence

Agata Pilitowska

Interval bilattices and some other simple bilattices

Wen-Fong Ke

Composition-diamond lemma for associative conformal algebras

12:15 - 12:45 Péter P. Pálfy

Black-box recognition of finite simple groups

AAA62 – Schedule of Lectures

Friday, June 15th, Afternoon

Lecture Hall HS 1 - “Raiffeisensaal”

Chair: Andrei Krokhin

14:30 - 15:20 Paweł M. Idziak

Generative Complexity in Algebra

Lecture Hall HS 3

Chair: Péter P. Pálfy

Lecture Hall HS 5

Chair: Vyachelav A. Artamonov

Lecture Hall HS 6

Chair: Shelly L. Wismath

Lecture Hall HS 7

Chair: Ivan Chajda

15:30 - 16:00 Andrei Kelarev

On complete and bipartite Cayley graphs

Petr Vojtechovsky

Octonions, finite simple Moufang loops, and groups of type G_{2}

Gerhard Dorfer

On congruences in orthomodular lattices

16:00 - 16:30 Peter Mayr

Fixed-point-free representations of finite groups and Frobenius Groups

Barbara

Roszkowska-Lech

Abelian Quandles

Sándor Radeleczki

The product decompositions of lattices

Miroslav Haviar

The “Full vs. Strong” problem in duality theory

16:30 - 17:00  Coffee Break

17:00 - 17:30 Evgeni Bashkirov

Irreducible linear groups of degree 3 over a quaternion division ring, containing a root subgroup

Eckehart Hotzel

Large right projective ideals in monoids with 0

Branimir Seselja

Complete meet irreducibles, closure operator and DM completions

Alex J. Lemin

On Gelfand’s problem concerning graphs, lattices, and ultrametric spaces

17:30 - 18:00 Ali Iranmanesh

A  characterization  of

PSU(3, q) for q > 5

Smile Markovski

On semigroups with identity xxy = y

John Rhodes

The q theory: Scott continuous functions on the algebraic lattice of all pseudovarieties of finite semigroups satisfying global Mal’cev conditions

Dragan Mašulović

Turning retractions of an algebra into an algebra

18:00 - 18:30 Peter Hubert

A construction of triply factorized groups using local near-rings

Jan Galuszka

Finite idempotent groupoids

Jietai Yu

Coordinates of Polynomial Algebras

19:00 - 22:00  Reception in Repräsentationsraum A (Uni-Center)

AAA62 – Schedule of Lectures

Saturday, June 16th, Morning

Lecture Hall HS 3

Chair: Kazimierz Głazek

Lecture Hall HS 4

Chair: Klaus Denecke

Lecture Hall HS 7

Chair: Dietmar Schweigert

9:00 - 9:50 E. Tamás Schmidt

Congruence preserving extensions of lattices

10:00 - 10:30 Howard E. Bell

Some rings related to periodic rings

Anna Romanowska

Varieties of binary modes

Jiří Tŭma

Amalgamation properties of lattices

10:30 - 11:00 Geoff Booth

Radicals whose semisimple classes satisfy a generalized ADS condition

Katarzyna Matczak

Quasivarieties of commutative binary modes

Géza Takách

Notes on pasting (joint work with Florence Micol)

11:00 - 11:30  Coffee break

11:30 - 12:00 Ladislav Bican

Relative purity over Noetherian rings

Anna Zamojska-Dzienio

Embedding sums of cancellative modes over a normal band into semimodules

Andreja Tepavčević

Special elements in posets

12:00 - 12:30 Rahul A. Manasi

A short note on the generation of a class of rings with subrings having different multiplicative identity than the ring

Dejan Delić

Badly structured classes of algebras

Miroslav Ploščica

Some properties of distributive congruence lattices

AAA62 – Schedule of Lectures

Saturday, June 16th, Afternoon

Lecture Hall HS 3

Chair: Ladislav Bican

Lecture Hall HS 5

Chair: Miroslav Haviar

Lecture Hall HS 7

Chair: Andreja Tepavčević

14:30 - 15:20 Kalle Kaarli

The equivalence lattices and the functions they support

15:30 - 16:00 Joanna Kaleta

Ternary representations of GBbi-QRs (Generalized Boolean bi-Quasirings)

Jens-Uwe Grabowski

Generalized relational varieties

Martin Goldstern

Clones on infinite sets: Results and open questions

16:00 - 16:30 Dietmar Dorninger

Ring-like structures with unique symmetric difference related to quantum logic

Grzegorz Binczak

Equational basies for monounary weak varieties

Jürgen Ecker

1-affine complete groups

16:30 - 17:00  Coffee break

17:00 - 17:30 Sergei V. Limarenko

Extension of density theorem for graded rings

Richard Holzer

Orthomodular partial algebras

Erhard Aichinger

The polynomial functions on certain semidirect products of groups

17:30 - 18:00 Mirjana Vukovic

About extra- and paragraded rings

Rudolf Wille

Restructuring General Algebra: An Approach Based on the Idea of Operational Descriptions

Nikita Artamonov

Spectral theory of differential operators

18:00 - 18:30 Kazimierz Głazek

Pseudo-nearrings and quasi-moduls over them

Gabriele Ricci

Two-ary integers can improve LISP-addressing and justify universal affine notions

AAA62 – Schedule of Lectures

Sunday, June 17th, Morning

Lecture Hall HS 4

Chair: Jens-Uwe Grabowski

Lecture Hall HS 5

Chair: Anna Romanowska

Lecture Hall HS 7

Chair: Rudolf Wille

9:00 - 9:50 Peter Hilton

The origins of homological algebra

10:00 - 10:30 Tim Boykett

Automata, Algebra, and Algorithms: Efficient Exhaustive Enumeration

Jiří Sichler

Coconnectedness

Ki-Bong Nam

Generalized Weyl algebra

10:30 - 11:00 Fred E. J. Linton

Associativity Problems for Tensor Products

Olga Sokratova

Onn pullback flatness of acts over Ω-rings

Vyachelav A. Artamonov

Quantum polynomials

11:00 - 11:30  Coffee Break

11:30 - 12:20 Carlton J. Maxson

Forcing Linearity Numbers

12:20 - 12:40 Reinhard Pöschel

AAA63 and AAA64